# Kinetically Modified Palatini Inflation Meets ACT Data

### v2

@@ Abstract
PACs numbers: 98.80.Cq, 04.50.Kd, 12.60.Jv, 04.65.+e
@@ II Set-up
Here $\mathfrak{g}$ is the determinant of the background Friedmann-Robertson-Walker metric, $g^{\mu\nu}$ with signature (+,,,)(+,-,-,-)( + , - , - , - ) and we allow for a kinetic mixing through the function $f_{\rm K}(\phi)$ which is conveniently parameterized in terms of $f_{\cal R}$ as follows
@@
is an indicative interval for $m$ which assures an ample variation of $f_{\rm K}$. Indeed, any value of $f_{\rm K}$ can be revealed by the parametrization of Eq. (6) selecting conveniently $m=\ln f_{\rm K}/\ln f_{\cal R}$. The considered $m$ margin can be characterized as natural, since the values are of order unity, and allows us to overlaid almost the whole observationally favored region, as we show in Sec. V. Contrary to the models in Ref. martin – with which ours share the same name – we do not violate the quadratic form of the kinetic operator ˙$\dot{}\phi^{2}$ for $\mu=\nu=0$. By performing a conformal transformation nmi , according to which we define the Einstein frame (EF) metric $\widehat{g}_{\mu\nu}=f_{\cal R}\,g_{\mu\nu}$, we can write ${\sf S}$ in the EF as follows
${\sf S}=\int d^{4}x\sqrt{-\widehat{\mathfrak{g}}}\left(-\frac{1}{2}\widehat{\mathcal{R}}+\frac{1}{2}\widehat{g}^{\mu\nu}\partial_{\mu}\widehat{\phi}\partial_{\nu}\widehat{\phi}-\widehat{V}_{\rm I}(\widehat{\phi})\right),$ (7)
@@ III Inflationary Dynamics
with the symbol $,\chi$ as subscript denoting derivation w.r.t the field $\chi$. For the expressions above we employ $J$ in Eq. (8), without expressing explicitly $\widehat{V}_{\rm I}$ in terms of $\widehat{\phi}$. For some $m$ values, actually, this is not doable due to the complicate form of the $\widehat{\phi}(\phi)$ function which is the following
@@
$a_{\rm s}$ $-(m+3)/(m+2)N^{2}_{\star}-{\cal O}\big{(}N^{-2}_{\star}(c_{\cal R}^{2}{N_{\star}})^{-1/(m+2)}\big{)}$ $-(m+2)/(m+1)N^{2}_{\star}-{\cal O}\big{(}N^{-2}_{\star}(c_{\cal R}{N_{\star}})^{-1/(m+1)}\big{)}$
@@
[…]
@@
[…]
@@ IV Effective Cut-Off Scale
$J^{2}\dot{\phi}^{2}=\dot{}\phi^{2}\begin{cases}1&\mbox{for}~{}~{}m=1,\\ 1+c_{\cal R}\phi^{n/2}&\mbox{for}~{}~{}m=2,\\ \left(1+2c_{\cal R}\phi^{n/2}+c_{\cal R}^{2}\phi^{n}+\cdots\right)&\mbox{for}~{}~{}m=3.\end{cases}$ (21)
@@
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for $n=2$ (left panel) and $4$ (right panel) and various $m$’s indicated on the curves. We take ${N_{\star}}=50$ (dashed lines) or ${N_{\star}}=60$ (solid lines) and minimal (thick lines) or maximal (thin lines) $c_{\cal R}$ values. The marginalized joint $68\%$ [$95\%$] c.l. regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
@@ V Numerical Results
$8\lesssim\frac{c_{\cal R}}{100}\lesssim 3.5\cdot 10^{3}$ or $4\displaystyle 9.7\lesssim\frac{c_{\cal R}}{100}\lesssim 4.2\cdot 10^{3}~{}~{}\mbox{for}~{}~{}n=$ 4 .
@@ VI Supergravity Embedding
${\sf S}=\int d^{4}x\sqrt{-\widehat{\mathfrak{g}}}\left(-\frac{1}{2}\widehat{\mathcal{R}}+K_{{\alpha}{\bar{\beta}}}\widehat{g}^{\mu\nu}\partial_{\mu}z^{\alpha}\partial_{\nu}z^{*{\bar{\beta}}}-\widehat{V}_{\rm F}\right)$ (29a)
@@
The form of $W$ can be uniquely determined if we impose an $R$ symmetry, under which $S$ and $\Phi$ have charges $1$ and 00, and a global $U(1)$ symmetry with assigned charges to $S$ and $\Phi$ $-1$ and $2/n$. The latter is violated, though, in the proposed $K$ which is judiciously chosen so that $J$ and $\widehat{V}_{\rm I}$ in Eq. (8) are reproduced. More specifically, $\widehat{V}_{\rm I}$ has to be derived from the only surviving term in Eq. (29b) along the track in Eq. (31), which is
@@
$1$ real $\widehat{\theta}$ $(6+3n^{2}c_{\cal R}/2\phi^{2-n/2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
$2$ real $\widehat{s},\widehat{\bar{s}}$ $(6/N_{2}+3n^{2}/2\phi^{2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
@@
$2$ Weyl $({\widehat{\psi}_{S}\pm\widehat{\psi}_{\Phi})/\sqrt{2}}~{}~{}$ $3n^{2}\widehat{H}_{\rm I}^{2}/2\phi^{2}f_{\cal R}^{m+1}$
@@
In particular, we find the expressions of the masses squared $\widehat{m}^{2}_{\chi^{\alpha}}$ (with $\chi^{\alpha}={\theta}$ and $s$) arranged in Table 2. These expressions assist us to appreciate the role of $N_{2}$ with $60<N_{2}<$ 6 in retaining positive $\widehat{m}^{2}_{s}$ – in practise we use $N_{2}=1$. Also we confirm that $\widehat{m}^{2}_{\chi^{\alpha}}\gg\widehat{H}_{\rm I}^{2}=\widehat{V}_{\rm I0}/3$ for $\phi_{\rm f}\leq\phi\leq\phi_{\star}$. In Table 2 we display the masses $\widehat{m}^{2}_{\psi^{\pm}}$ of the corresponding fermions too. We define $\widehat{\psi}_{S}=\sqrt{K_{SS^{*}}}\psi_{S}$ and $\widehat{\psi}_{\Phi}=\sqrt{K_{\Phi\Phi^{*}}}\psi_{\Phi}$ where $\psi_{\Phi}$ and $\psi_{S}$ are the Weyl spinors associated with $S$ and $\Phi$ respectively.
@@ VII Conclusions
Namely, working along the lines of Ref. nmikr , we considered a non-canonical kinetic mixing, Eq. (6), – involving the exponent $m$ – in the inflaton sector, apart from the non-minimal coupling to gravity, Eq. (1), which is associated with the potential in Eq. (2). Confining $m$ to the range $(0.87~{}[0.5]-10)$ for $n=2$ [$n=4$] we achieved observational predictions which offer a nice covering of the present data on $n_{\rm s}$ – its compatibility especially with the $n=4$ case for low $m$ values is really impressive, as shown in Fig. 1. Part of the resulting $r$ values there could be accessible in the near future. Our solutions can be attained with subplanckian values of the inflaton, requiring large $c_{\cal R}$’s, and without causing serious problems with the perturbative unitarity since the unitarity-violation scale remains well above the inflationary one. It is gratifying, in addition, that a sizable fraction of the allowed parameter space of our models (with $c_{\cal R}\gg 1$) can be studied analytically.

### v3

@@ Abstract
PACs numbers: 98.80.Cq, 04.50.Kd, 12.60.Jv, 04.65.+e Published in Phys. Lett. B 868, 139739 (2025)
@@ II Set-up
Here $\mathfrak{g}$ is the determinant of the background Friedmann-Robertson-Walker metric, $g^{\mu\nu}$ with signature (+,,,)(+,-,-,-) and we allow for a kinetic mixing through the function $f_{\rm K}(\phi)$ which is conveniently parameterized in terms of $f_{\cal R}$ as follows
@@
is an indicative interval for $m$ which assures an ample variation of $f_{\rm K}$. Indeed, any value of $f_{\rm K}$ can be revealed by the parametrization of Eq. (6) selecting conveniently $m=\ln f_{\rm K}/\ln f_{\cal R}$. The considered $m$ margin can be characterized as natural, since the values are of order unity, and allows us to overlaid almost the whole observationally favored region, as we show in Sec. V. Contrary to the models in Ref. martin – with which ours share the same name – we do not violate the quadratic form of the kinetic operator $\dot{}\phi^{2}$ for $\mu=\nu=0$. By performing a conformal transformation nmi , according to which we define the Einstein frame (EF) metric $\widehat{g}_{\mu\nu}=f_{\cal R}\,g_{\mu\nu}$, we can write $\sf S$} in the EF as follows
[…]
@@ III Inflationary Dynamics
with the symbol $\chi$ as subscript denoting derivation w.r.t the field $\chi$. For the expressions above we employ $J$ in Eq. (8), without expressing explicitly $\widehat{V}_{\rm I}$ in terms of $\widehat{\phi}$. For some $m$ values, actually, this is not doable due to the complicate form of the $\widehat{\phi}(\phi)$ function which is the following
@@
$a_{\rm s}$ $(m+3)/(m+2)N^{2}_{\star}-{\cal O}\big{(}N^{-2}_{\star}(c_{\cal R}^{2}{N_{\star}})^{-1/(m+2)}\big{)}$ $(m+2)/(m+1)N^{2}_{\star}-{\cal O}\big{(}N^{-2}_{\star}(c_{\cal R}{N_{\star}})^{-1/(m+1)}\big{)}$
@@
${N_{\star}}=\int_{\widehat{\phi}_{\rm f}}^{\widehat{\phi}_{\star}}d\widehat{\phi}\frac{\widehat{V}_{\rm I}}{\widehat{V}_{\rm I,\widehat{\phi}}}~~\mbox{and}~~A_{\rm s}^{1/2}=\frac{1}{2\sqrt{3}\,\pi}\;\frac{\widehat{V}_{\rm I}^{3/2}(\widehat{\phi}_{\star})}{|\widehat{V}_{\rm I,\widehat{\phi}}(\widehat{\phi}_{\star})|}$, (13)
@@
$n_{\rm s}=\>1-6\widehat{\epsilon}_{\star}\ +\ 2\widehat{\eta}_{\star},~~r=16\widehat{\epsilon}_{\star}$, (18a)
$a_{\rm s}=\>2\left(4\widehat{\eta}_{\star}^{2}-(n_{\rm s}-1)^{2}\right)/3-2\widehat{\xi}_{\star}$, (18b)
@@ IV Effective Cut-Off Scale
$J^{2}\dot{\phi}^{2}=\dot{}\phi^{2}\begin{cases}1&\mbox{for}~~m=1,\\ 1+c_{\cal R}\phi^{n/2}&\mbox{for}~~m=2,\\ \left(1+2c_{\cal R}\phi^{n/2}+c_{\cal R}^{2}\phi^{n}+\cdots\right)&\mbox{for}~~m=3.\end{cases}$
@@
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for $n=2$ (left panel) and $4$ (right panel) and various $m$’s indicated on the curves. We take ${N_{\star}}=50$ (dashed lines) or ${N_{\star}}=60$ (solid lines) and minimal (thick lines) or maximal (thin lines) $c_{\cal R}$ values. The marginalized joint $68\%$ $95\%$] c.l. regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
@@ V Numerical Results
$8\lesssim\frac{c_{\cal R}}{100}\lesssim 3.5\cdot 10^{3}$ or $9.7\lesssim\frac{c_{\cal R}}{100}\lesssim 4.2\cdot 10^{3}~~\mbox{for}~~n=4.$
@@ VI Supergravity Embedding
$\hskip-5.69054pt{\sf S}=\int d^{4}x\sqrt{-\widehat{\mathfrak{g}}}\left(-\frac{1}{2}\widehat{\mathcal{R}}+K_{{\alpha}{\bar{\beta}}}\widehat{g}^{\mu\nu}\partial_{\mu}z^{\alpha}\partial_{\nu}z^{*{\bar{\beta}}}-\widehat{V}_{\rm F}\right)$ (28a)
@@
The form of $W$ can be uniquely determined if we impose an $R$ symmetry, under which $S$ and $\Phi$ have charges $1$ and 0, and a global $U(1)$ symmetry with assigned charges to $S$ and $\Phi$ 1-1 and $2/n$. The latter is violated, though, in the proposed $K$ which is judiciously chosen so that $J$ and $\widehat{V}_{\rm I}$ in Eq. (8) are reproduced. More specifically, $\widehat{V}_{\rm I}$ has to be derived from the only surviving term in Eq. (28b) along the track in Eq. (30), which is
@@
11 real $\widehat{\theta}$ $(6+3n^{2}c_{\cal R}/2\phi^{2-n/2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
22 real $\widehat{s},\widehat{\bar{s}}$ $(6/N_{2}+3n^{2}/2\phi^{2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
@@
22 Weyl $({\widehat{\psi}_{S}\pm\widehat{\psi}_{\Phi})/\sqrt{2}}$~~ $3n^{2}\widehat{H}_{\rm I}^{2}/2\phi^{2}f_{\cal R}^{m+1}$
@@
In particular, we find the expressions of the masses squared $\widehat{m}^{2}_{\chi^{\alpha}}$ (with $\chi^{\alpha}={\theta}$ and $s$) arranged in Table 2. These expressions assist us to appreciate the role of $N_{2}$ with $0<N_{2}<6$ in retaining positive $\widehat{m}^{2}_{s}$ – in practise we use $N_{2}=1$. Also we confirm that $\widehat{m}^{2}_{\chi^{\alpha}}\gg\widehat{H}_{\rm I}^{2}=\widehat{V}_{\rm I0}/3$ for $\phi_{\rm f}\leq\phi\leq\phi_{\star}$. In Table 2 we display the masses $\widehat{m}^{2}_{\psi^{\pm}}$ of the corresponding fermions too. We define $\widehat{\psi}_{S}=\sqrt{K_{SS^{*}}}\psi_{S}$ and $\widehat{\psi}_{\Phi}=\sqrt{K_{\Phi\Phi^{*}}}\psi_{\Phi}$ where $\psi_{\Phi}$ and $\psi_{S}$ are the Weyl spinors associated with $S$ and $\Phi$ respectively.
@@ VII Conclusions
Namely, working along the lines of Ref. nmikr , we considered a non-canonical kinetic mixing, Eq. (6), – involving the exponent $m$ – in the inflaton sector, apart from the non-minimal coupling to gravity, Eq. (1), which is associated with the potential in Eq. (2). Confining $m$ to the range (0.87[0.5]10)(0.87~[0.5]-10) for $n=2$ [$n=4$] we achieved observational predictions which offer a nice covering of the present data on $n_{\rm s}$ – its compatibility especially with the $n=4$ case for low $m$ values is really impressive, as shown in Fig. 1. Part of the resulting $r$ values there could be accessible in the near future. Our solutions can be attained with subplanckian values of the inflaton, requiring large $c_{\cal R}$’s, and without causing serious problems with the perturbative unitarity since the unitarity-violation scale remains well above the inflationary one. It is gratifying, in addition, that a sizable fraction of the allowed parameter space of our models (with $c_{\cal R}\gg 1$) can be studied analytically.